# Lattice-induced sound trapping in biperiodic metasurfaces of acoustic resonators

### v3

@@ III.3 Realization in the infinite lattice
Thus, an infinite biperiodic lattice of spherical resonators can support the acoustic BIC due to the interference of the zonal multipole of a certain parity [see Fig. 1(b)]. The case of even parity has been discussed in this section, whereas the odd-parity BIC is considered in Appendix I. Finally, let us also discuss the impact of the lattice geometry. The form of matrix (25) will be the same for the hexagonal lattice $\mathfrak{n}=6$ in (12)] but with different values of the lattice sums. The monoclinic and orthorhombic lattices ($\mathfrak{n}=2$) allow the coupling of the zonal multipoles $a_{0,0}^{\rm eff}$, and $a_{2,0}^{\rm eff}$] and the nonzonal quadrupoles $a_{2,\pm 2}^{\rm eff}$, which however do not radiate acoustic waves in the normal direction. For arbitrarily shaped scatterers, the off-diagonal entries of the single-particle T-matrix can also be nonzero. In this case, Eq. (25) has to consider them as well.
@@ IV Quasi-BIC regime
As shown in the previous section, infinite lattices of spherical particles can support acoustic BICs resulting from multipole interference at specific frequencies and lattice constants. By tuning the lattice constant, the BIC with an infinite $Q$-factor can be transformed into a quasi-BIC, which manifests itself in the spectrum as a Fano resonance with a large $Q$-factor compared to resonances of an isolated resonator. In practice, the quasi-BIC can be more important in applications than the genuine BIC, since it can be excited by a normally incident plane wave. The acoustic quasi-BICs have already been utilized to enhance the acoustic Purcell factor [75, 13], and to achieve perfect absorption [23, 76] or localization of sound energy [16]. In this section, we discuss the acoustic quasi-BIC regime under realistic conditions, where either material losses or a substrate is present, and also its manifestation in finite-size arrays.
The total $Q$-factor of an eigenmode can be decomposed into two terms as
@@
where $Q_{\rm rad}$ and $Q_{\rm nr}$ are responsible for radiative and nonradiative (material) losses, respectively. For a genuine BIC in the infinite lattice, $Q_{\rm rad}^{-1}=0$ and consequently the $Q$-factor is limited by $Q_{\rm tot}=Q_{\rm nr}$ even for a genuine BIC [12]. For a quasi-BIC in both finite and infinite structures, $Q_{\rm rad}^{-1}>0$ and Eq. (27) shall be used to estimate the upper bound of the $Q$-factor.
@@ IV.1 Influence of material absorption and substrate
First, we assume that the particles have material losses by considering the complex-valued acoustic impedance $Z_{\rm s}=Z_{\rm s}^{\prime}+\mathrm{i}Z_{\rm s}^{\prime\prime}$ where $Z_{\rm s}^{\prime}=\rho_{\rm s}c_{\rm s}=0.216$ MRayl. The ratio of the imaginary part to the real part defines the so-called loss tangent $\tan\delta=Z_{\rm s}^{\prime\prime}/Z_{\rm s}^{\prime}$. Figure 4(a) shows the transmittance spectra of the lattice with the normalized lattice constant of $L/a=1.5$ for different values of the loss tangent, while Fig. 4(b) shows the corresponding absorptance spectra. Fig. 4(c) plots the values of the total $Q$-factor of the quasi-BIC, which are estimated using the procedure from Ref. 73 to fit the resonances in Fig. 4(a). For the resonators without material losses, the quasi-BIC has the radiative $Q$-factor and consequently the total $Q$-factor of $Q_{\rm tot}=Q_{\rm rad}\approx 129$, whereas the absorptance is zero [see the blue curves in Figs. 4(a) and 4(b)]. Meanwhile, one can observe that the larger the material losses, the lower the total $Q$-factor in Fig. 4(c) and the larger absorptance at the frequency of the quasi-BIC in Fig. 4(b), while the resonance in the transmittance spectrum in Fig. 4(a) weakens. Thus, the acoustic quasi-BIC resonance in a lossy system appears as an absorptance resonance, as well as its electromagnetic counterpart [28]. Moreover, for the lossy resonators, the decrease in the nonradiative $Q$-factor with the loss tangent can be justified by the approximate expression on the upper bound of the $Q$-factor given by $Q_{\rm nr}=1/\tan\delta$ [73].
@@
Figure 4: Influence of material losses in scatterers (resonators) on the quasi-BIC resonance. (a) Transmittance of the lattice $|t|^{2}$ [see Eq. (16b)] in the quasi-BIC regime for different values of material loss tangent being 0 (blue), $10^{-3}$ (orange), and $10^{-2}$ (green). The normalized lattice constant is $L/a=1.5$. (b) The corresponding absorptance $(1-|r|^{2}-|t|^{2})$ of the lattice. (c) Total $Q$-factor of the quasi-BIC as a function of material loss tangent.
@@
Similarly, we can analyze the impact of a substrate. Since the considered BIC emerges for the specific values of the parameters, it is sensitive to symmetry breaking, as well as other accidental BICs [12]. Hence, the larger the impedance contrast between the upper and lower half-spaces, the smaller the $Q$-factor. Figure 5(a) shows the transmittance spectra of the lattice with the normalized lattice constant of $L/a=1.5$ for different substrates. The material parameters of the substrates are taken from Ref. 53. Figure 5(b) plots the $Q$-factor of the resonance as a function of the normalized substrate impedance $Z_{\rm sub}/Z_{\rm b}$ where $Z_{\rm b}=\rho_{\rm b}c_{\rm b}=1.497$ MRayl is the impedance of the upper half-space (note that $Q_{\rm nr}^{-1}=0$). One can clearly see that if the impedance contrast $Z_{\rm sub}/Z_{\rm b}$ increases, the $Q$-factor decreases. Thus, acoustic metasurfaces in the quasi-BIC regime are strongly responsive to background variations and can therefore serve as the basis for highly sensitive acoustic sensors.
@@
Figure 5: Influence of a substrate on the quasi-BIC resonance. (a) Transmittance of the lattice $|t|^{2}$ [see Eq. (16b)] in the quasi-BIC regime for different substrates. The normalized lattice constant is $L/a=1.5$. (b) $Q$-factor of the quasi-BIC as a function of the normalized substrate impedance $Z_{\rm sub}/Z_{\rm b}$ with $Z_{\rm b}=1.497$ MRayl.
IV.2 Realization in the finite-size arrays
@@
The considered BIC has an infinite $Q$-factor because the lattice and eigenmode symmetry allow for only one direction of radiation, being the $z$ axis. In contrast, the scattering amplitudes (19) for the directions z-z and +z+z are adjusted to zero at the BIC resonance due to (26). Therefore, the overall radiation of the lattice eigenmode is zero and the mode is trapped (BIC). In the experiment, all structures have finite sizes along all dimensions. Since a finite-size array has an infinite number of allowed radiation channels, it always possesses nonzero radiation losses [77, 10, 12]. Thus, finite-size structures cannot possess either electromagnetic or acoustic BICs with an infinite $Q$-factor. Note that elastic particles, which allow for shear waves, in a nonelastic and nonviscous medium, can possess a genuine BIC [22], but we do not consider them in this work. Thus, if a lattice of acoustic particles becomes a finite-size $N\times N$ array, the considered BIC always transforms into a quasi-BIC. Moreover, the response of finite-size arrays can remarkably differ from that of an infinite lattice [78, 79, 80]. Hence, we also discuss the development of the BIC resonance for finite arrays.
@@
To simulate the acoustic response of finite-size arrays, we also use the T-matrix-based formalism. Within the approach, the scattered pressure field of the $i$-th sphere is again described by a vector of multipole coefficients $\mathbf{a}(\mathbf{r}_{i})$ as well as in the infinite lattice case. However, because of the broken translation symmetry, the coefficients are no longer linked through the Bloch theorem, but via a system of linear equations considering the interparticle coupling (see Appendix J). A solution to system (69), which can be obtained using acoustotreams code, provides us with the self-consistent response of the finite-size array. We can also calculate the scattering cross section of the $N\times N$ array $\sigma_{N}(\omega)$ and divide it by $N^{2}\sigma_{\rm geom}$ to obtain the effective scattering efficiency of a particle in the array. In Fig. 6, it is plotted as a function of frequency and $N$ when the normalized lattice constant is fixed at the value of $L/a=1.5$ corresponding to the quasi-BIC resonance in the infinite lattice. First, one can notice that even the response of the $3\times 3$ array differs significantly from that of the single scatterer. Namely, the quasi-BIC resonance emerges already for $N=3$ that corresponds to a minimum of the scattering efficiency. Moreover, as $N$ increases, the resonance frequency approaches the limit of the infinite lattice and converges for $N=11$.
Figure 6: Effective scattering efficiency per resonator for an $N\times N$ array of spherical resonators with the normalized lattice constant of $L/a=1.5$. The red line indicates the frequency of the quasi-BIC resonance of the infinite lattice with the same lattice constant.
@@ V Conclusion
In this work, we investigate acoustic bound states in the continuum (trapped modes) in two-dimensional metasurfaces of scatterers (acoustic resonators), with one resonator per unit cell and possessing inversion symmetry. First, we develop a comprehensive theoretical model that enables us to obtain the analytical conditions of the BIC occurrence. Next, we show that an acoustic BIC can emerge at the $\Gamma$-point of the Brillouin zone due to the interference of effective zonal multipoles of certain parity within a metasurface possessing inversion symmetry. Our numerical simulations reveal a spectral feature associated with the BIC in the acoustic spectrum of the lattice under normal incidence. Furthermore, a slight detuning of the lattice constant transforms the genuine BIC into a quasi-BIC, which manifests in the spectrum as a Fano resonance.
Focusing on the experiment, we also consider the quasi-BIC regime under realistic assumptions including material losses in the resonators, and a substrate. Moreover, we investigate the development of the quasi-BIC resonance in finite $N\times N$ arrays of particles as a function of $N$ and identify the values of $N$ sufficient for convergence to the infinite lattice limit.
Our results not only provide the theoretical and numerical foundations for acoustic BICs in metasurfaces with inversion symmetry but also demonstrate their practical feasibility under realistic conditions. The ability to control the transition from genuine BICs to quasi-BICs through simple geometric tuning provides a versatile mechanism for engineering highly selective, low-loss acoustic responses. This approach opens new opportunities for designing advanced acoustic metasurfaces with tunable functionalities. Potential applications range from ultra-narrowband acoustic filters and high-sensitivity sensors to energy-harvesting devices and medical ultrasound technologies. By bridging fundamental physics and practical design strategies, our work paves the way for next-generation acoustic systems that exploit the unique properties of BICs for enhanced wave control and device performance.
@@ Appendix A Scalar spherical waves and their translation coefficients
The scalar spherical multipoles obey the following addition theorem [82]
$\Psi^{(3)}_{\ell,m}(k;\mathbf{r})=\sum_{\ell^{\prime},m^{\prime}}\mathcal{A}^{(3)}_{\ell^{\prime},m^{\prime},\ell,m}(k;\mathbf{r}-\mathbf{r}_{0})\Psi^{(1)}_{\ell^{\prime},m^{\prime}}(k;\mathbf{r}_{0})\,.$ (31)
The translation coefficients of (28) are given by
@@
Appendix B Expansion of a plane wave into scalar spherical multipoles
A pressure plane wave $p_{\mathrm{inc}}(\omega;\mathbf{r})=p_{0}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}$ can be expanded into regular scalar spherical waves (28) as follows
@@ Appendix B Expansion of a plane wave into scalar spherical multipoles
where $\theta_{\mathbf{k}}$ and $\varphi_{\mathbf{k}}$ are the spherical angles of $\mathbf{k}$ with $|\mathbf{k}|=\omega/c$. Hence, the expansion coefficients of the plane wave are $b_{l,m}=p_{0}4\pi\mathrm{i}^{\ell}Y^{*}_{\ell,m}(\theta_{\mathbf{k}},\varphi_{\mathbf{k}})$.

### v4

@@ III.3 Realization in an infinite metasurface
Furthermore, the resonance associated with the BIC can also be observed in the transmittance spectrum for a normally incident pressure plane wave $p_{\rm inc}(\mathbf{r})=\mathrm{e}^{\mathrm{i}kz}$. To find the metasurface response, we calculate effective multipole moments (11) up to $\ell_{\rm max}=6$ and then calculate transmission (16b) using acoustotreams. Figure 3(b) displays the transmittance $T\equiv|t|^{2}$ of the metasurface as a function of the normalized lattice constant and size parameter, while Figs. 3(e)-(g) plot it as a function of the size parameter for the different values of the normalized lattice constant. In Fig. 3(b), one can see that the transmittance exhibits a resonant feature associated with the BIC for the same lattice constant and frequency where the lowest singular value vanishes. Note that exactly at the frequency of the BIC formed by the resonant monopole and quadrupole, the resonance disappears from the spectrum because the incident wave cannot excite a nonradiant mode, and the transmittance is therefore entirely determined by the nonresonant multipoles (dipole and octupole) [see also Fig. 3(f)]. However, as shown in Figs. 3(e) and 3(g), one can observe the development of Fano-type resonances (quasi-BICs) with relatively high quality factors ($Q$ factors) by gradually varying the lattice constant. Note that the Fano profile of the considered resonances arises from the interference (overlapping) of the incident field and the fields of the even multipoles (monopole and quadrupole) from the resonant quasi-BIC mode, with the addition of odd multipoles (dipole and octupole) from the nonresonant mode [29]. To estimate the $Q$-factor of a Fano-type profile of the quasi-BIC, we fit the transmittance spectrum for a given lattice constant with a sum of several Fano resonances as described in Ref. 77. In Figure 3(h), the $Q$-factor decreases as the inverse square of the normalized lattice constant offset $|\Delta L|/a\equiv|L-L_{\rm BIC}|/a$, where the normalized lattice constant of the BIC is $L_{\rm BIC}/a=1.395$, since Eq. (26) ceases to hold when $|\Delta L|/a>0$. Note that the $Q$-factor of the genuine BIC at $L_{\rm BIC}/a$ is infinite. Finally, to verify the transmittance calculated with the T-matrix method, we also compare it in Figs. 3(e)-(g) with that computed by the pressure acoustics module of COMSOL Multiphysics™, v.5.5 [78]. Appendix K concisely describes the model. The figures confirm that both methods demonstrate perfect agreement.
@@ IV Quasi-BIC regime
As shown in the previous section, infinite lattices of spherical particles can support acoustic BICs resulting from multipole interference at specific frequencies and lattice constants. By tuning the lattice constant, the BIC with an infinite $Q$ factor can be transformed into a quasi-BIC, which manifests itself in the spectrum as a Fano resonance with a large $Q$ factor compared to resonances of an isolated resonator. In practice, the quasi-BIC can be more important in applications than the genuine BIC, since it can be excited by a normally incident plane wave. The acoustic quasi-BICs have already been utilized to enhance the acoustic Purcell factor [79, 13], and to achieve perfect absorption [23, 80] or localization of sound energy [16]. In this section, we discuss the acoustic quasi-BIC regime under realistic conditions, where either material losses or a substrate is present, and also its manifestation in finite-size arrays.
The total $Q$ factor of an eigenmode can be decomposed into two terms as
@@
where $Q_{\rm rad}$ and $Q_{\rm nr}$ are responsible for radiative and nonradiative (material) losses, respectively. For a genuine BIC in the infinite lattice, $Q_{\rm rad}^{-1}=0$ and consequently the $Q$ factor is limited by $Q_{\rm tot}=Q_{\rm nr}$ even for a genuine BIC [12]. For a quasi-BIC in both finite and infinite structures, $Q_{\rm rad}^{-1}>0$ and Eq. (28) should be used to calculate the $Q$ factor.
@@ IV.1 Influence of material absorption and substrate
First, we assume that the particles have material losses by considering the complex-valued acoustic impedance $Z_{\rm s}=Z_{\rm s}^{\prime}+\mathrm{i}Z_{\rm s}^{\prime\prime}$, where $Z_{\rm s}^{\prime}=\rho_{\rm s}c_{\rm s}=0.216$ MRayl. The ratio of the imaginary part to the real part defines the so-called loss tangent $\tan\delta=Z_{\rm s}^{\prime\prime}/Z_{\rm s}^{\prime}$. Figure 4(a) shows the transmittance spectra of the metasurface with the normalized lattice constant of $L/a=1.5$ for different values of the loss tangent, while Fig. 4(b) shows the corresponding absorptance spectra. Figure 4(c) plots the values of the total $Q$ factor of the quasi-BIC, which are estimated using the procedure from Ref. 77 to fit the resonances in Fig. 4(a). For the resonators without material losses, the quasi-BIC has the radiative $Q$ factor and consequently the total $Q$ factor of $Q_{\rm tot}=Q_{\rm rad}\approx 129$, whereas the absorptance is zero [see the blue curves in Figs. 4(a) and 4(b)]. Meanwhile, one can observe that, as the material losses increase, the total $Q$ factor decreases [Fig. 4(c)], the absorptance at the quasi-BIC frequency increases [Fig. 4(b)], and the resonance in the transmittance spectrum weakens [Fig. 4(a)]. Thus, an acoustic quasi-BIC resonance in a lossy system appears as an absorptance resonance, as well as its electromagnetic counterpart [29]. Moreover, for the lossy resonators, the decrease in the nonradiative $Q$ factor with increasing loss tangent can be justified by its approximate upper-bound expression, $Q_{\rm nr}=1/\tan\delta$ [77].
@@
Figure 4: Influence of material losses in scatterers (resonators) on the quasi-BIC resonance. (a) Transmittance of the metasurface $|t|^{2}$ [see Eq. (16b)] in the quasi-BIC regime for different values of material loss tangent being 0 (blue), $10^{-3}$ (orange), and $10^{-2}$ (green). The normalized lattice constant is $L/a=1.5$. (b) The corresponding absorptance $(1-|r|^{2}-|t|^{2})$ of the metasurface. (c) Total $Q$ factor of the quasi-BIC as a function of material loss tangent.
@@
Similarly, we can analyze the impact of a substrate. Since the considered BIC emerges for the specific values of the parameters, it is sensitive to symmetry breaking, as well as other accidental BICs [12]. Hence, the larger the impedance contrast between the upper and lower half-spaces, the smaller the $Q$ factor. Figure 5(a) shows the transmittance spectra of the metasurface with the normalized lattice constant of $L/a=1.5$ for different substrates. The material parameters of the substrates are taken from Ref. 54. Figure 5(b) plots the $Q$ factor of the resonance as a function of the normalized substrate impedance $Z_{\rm sub}/Z_{\rm b}$, where $Z_{\rm b}=\rho_{\rm b}c_{\rm b}=1.497$ MRayl is the impedance of the upper half-space (note that $Q_{\rm nr}^{-1}=0$). One can clearly see that if the impedance contrast $Z_{\rm sub}/Z_{\rm b}$ increases, the $Q$ factor decreases. Thus, acoustic metasurfaces in the quasi-BIC regime are strongly responsive to background variations and can therefore serve as the basis for highly sensitive acoustic sensors.
@@
Figure 5: Influence of a substrate on the quasi-BIC resonance. (a) Transmittance of the metasurface $|t|^{2}$ [see Eq. (16b)] in the quasi-BIC regime for different substrates. The normalized lattice constant is $L/a=1.5$. (b) $Q$ factor of the quasi-BIC as a function of the normalized substrate impedance $Z_{\rm sub}/Z_{\rm b}$ with $Z_{\rm b}=1.497$ MRayl.
IV.2 Realization in finite-size arrays
@@
The considered BIC has an infinite $Q$ factor because translation symmetry of the lattice allows for only one radiation channel, being the $z$ axis, whereas the scattering amplitudes (19) for the directions z-z and +z+z are zero for the BIC due to Eq. (26). Therefore, the overall radiation of the lattice eigenmode is zero, and the mode is trapped (BIC). In the experiment, all structures have finite sizes along all dimensions. Since a finite-size array has an infinite number of allowed radiation channels, it always possesses nonzero radiation losses [81, 10, 12]. Thus, finite-size structures can possess neither electromagnetic nor acoustic BICs with an infinite $Q$ factor. Note that elastic spherical resonators, which support shear waves, can possess a genuine transversely polarized BIC due to the polarization orthogonality between the fields inside and outside the resonator [22], but we do not consider them in this work. Thus, if a lattice of acoustic resonators becomes a finite-size $N\times N$ array, the considered BIC always transforms into a quasi-BIC. Moreover, the response of finite-size arrays can remarkably differ from that of an infinite lattice [82, 83, 84]. Hence, we also discuss the development of the BIC resonance for finite arrays.
@@
To simulate the acoustic response of finite-size arrays, we also use the T-matrix-based formalism. Within the approach, the scattered pressure field of the $i$th sphere is again described by a vector of multipole coefficients $\mathbf{a}(\mathbf{r}_{i})$ as well as in the infinite-lattice case. However, because of the broken translation symmetry, the coefficients are no longer linked through the Bloch theorem, but via a system of linear equations considering mutual coupling between resonators (see Appendix J). A solution to system (69), which can be obtained using acoustotreams code, provides us with the self-consistent response of the finite-size array. We can also calculate the scattering cross section of the $N\times N$ array $\sigma_{N}(\omega)$ and divide it by $N^{2}\sigma_{\rm geom}$ to obtain the effective scattering efficiency of a particle in the array. In Fig. 6(a), it is plotted as a function of frequency and $N$ when the normalized lattice constant is fixed at the value of $L/a=1.5$ corresponding to the quasi-BIC resonance in the infinite metasurface [see Fig. 3(e)]. First, it is notable that even the response of the $3\times 3$ array differs significantly from that of the single scatterer. Namely, the quasi-BIC resonance appears already for $N=3$, corresponding to a minimum in the scattering efficiency. Moreover, as $N$ increases, the resonance frequency approaches the infinite-array limit and converges by $N=11$.
Figures 6(b) and 6(c) show the multipole moments of the resonators on the $11\times 11$ array and the pressure field, generated by this array in the near-field region. One can see that the farther a resonator is from the center, the lower the near field is above it and the weaker the multipole moments are induced in the resonator, forming approximately a standing-wave pattern between the edges of the array with the wave vector $k_{x}=k_{y}=\frac{\pi}{(N+1)L}$ [see Eq. (2) in Ref. 85]. Thus, a normally incident plane wave mostly excites resonators that are close to the center, yielding strong (enhanced) near fields. In particular, it implies that the quasi-BIC should be better coupled to an external source that is located above the array center rather than above its edges.
[…]
@@ V Conclusion
In this work, we investigate acoustic bound states in the continuum (trapped modes) in biperiodic metasurfaces with one acoustic resonator (scatterer) per unit cell. First, we develop a comprehensive theoretical model based on the T-matrix method and symmetry analysis that enables us to obtain the analytical conditions of the BIC formation. Next, we show that an accidental acoustic BIC can emerge at the $\Gamma$-point of a metasurface with inversion symmetry due to destructive interference between two effective zonal multipoles of the same parity. Our numerical simulations reveal a spectral feature associated with the BIC in the spectrum of the metasurface when a pressure plane wave is normally incident. Furthermore, a slight detuning of the lattice constant transforms the genuine BIC into a quasi-BIC, which manifests in the spectrum as a Fano resonance.
Focusing on the experiment, we also consider the quasi-BIC regime under realistic assumptions, including either material losses in the resonators or a substrate. Moreover, we investigate the development of the quasi-BIC resonance in finite $N\times N$ arrays of particles as a function of $N$ and identify the values of $N$ sufficient for convergence to the infinite-lattice limit.
Our results not only provide the theoretical and numerical foundations for acoustic BICs in metasurfaces with inversion symmetry but also demonstrate their practical feasibility under realistic conditions. The ability to control the transition from genuine BICs to quasi-BICs through simple geometric tuning provides a versatile mechanism for engineering highly selective, low-loss acoustic responses. This approach opens opportunities for designing advanced acoustic metasurfaces with tunable functionalities. Potential applications range from ultranarrowband acoustic filters and high-sensitivity sensors to energy-harvesting devices and medical ultrasound technologies. By bridging fundamental physics and practical design strategies, our work paves the way for next-generation acoustic systems that exploit the unique properties of BICs for enhanced wave control and device performance.
@@ Appendix A Scalar spherical waves and their translation coefficients
The scalar spherical multipoles obey the following addition theorem [87]:
$\Psi^{(3)}_{\ell,m}(\omega;\mathbf{r})=\sum_{\ell^{\prime},m^{\prime}}\mathcal{C}^{(3)}_{\ell^{\prime},m^{\prime},\ell,m}(\omega;\mathbf{r}-\mathbf{r}_{0})\Psi^{(1)}_{\ell^{\prime},m^{\prime}}(\omega;\mathbf{r}_{0})\,.$ (32)
The translation coefficients of Eq. (29) are given by
@@
Appendix B Expansion of a plane wave into scalar spherical waves
A pressure plane wave $p_{\mathrm{inc}}(\omega;\mathbf{r})=p_{0}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}$ can be expanded into regular scalar spherical waves (29) as follows:
@@ Appendix B Expansion of a plane wave into scalar spherical waves
where $\theta_{\mathbf{k}}$ and $\varphi_{\mathbf{k}}$ are the spherical angles of $\mathbf{k}$ with $|\mathbf{k}|=\omega/c$. Hence, the expansion coefficients of the plane wave are $b_{\ell,m}=p_{0}4\pi\mathrm{i}^{\ell}Y^{*}_{\ell,m}(\theta_{\mathbf{k}},\varphi_{\mathbf{k}})$.